OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n3 + 2n is divisible by 3 for every positive integer n. PLEASE HELP I GIVE MEDALS PLEASE HELP

OpenStudy (anonymous):

1. for n=1 statement is true, we can check it: 1+2=3 => 3 divides 3 2. Assume that the statement is true for n=k: 3 divides k^3+2k 3. Then it has to be right for n=k+1 - let's check (k+1)^3+2(k+1)=k^3+3k^2+5k+3=(k^3+2k) + (3k^2+3k+3) = (k^3+2k) + 3(k^2+k+1) - we can divide this by 3, so the statement is proved

OpenStudy (anonymous):

Is that how id answer the question fully? @wintry

OpenStudy (anonymous):

how would I answer this one as it is?

OpenStudy (anonymous):

I don't know your requirements, this is how i was taught to solve this type of problems...

OpenStudy (anonymous):

I dont get what to do in the question ? @wintry

OpenStudy (anonymous):

Do you know what is math. induction?

OpenStudy (anonymous):

I got it .... can you help with: A certain species of tree grows an average of 3.8 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 5 meters tall. Help explaining

OpenStudy (anonymous):

@wintry

OpenStudy (anonymous):

Well, it's easy one. H(n) - is the sequence of heights of the tree every week, n is the number of weeks H(n)=5+0.038*n

OpenStudy (anonymous):

Sorry, it's in meters I'll rewrite it in cm: H(n)=500+3.8*n

OpenStudy (anonymous):

why 500 @wintry

OpenStudy (anonymous):

because 5 meters = 500 cm, you cannot add meters and cm, you should work in only system of measurements (I'm not sure whether I explain it correctly, English is not my native language)

OpenStudy (anonymous):

you explain fine

OpenStudy (anonymous):

is H(n)=500+3.8*n the answer or is there a different answer @wintry

OpenStudy (anonymous):

It has to be the answer, if I've got the problem correctly

OpenStudy (anonymous):

I also have trouble with these two @wintry

OpenStudy (anonymous):

just the options I thought the second one was B but Im not sure

OpenStudy (anonymous):

B is in the first, and I'm not sure about the second one. It has to be A, but it might be D cause I'm not sure whether this sequence ends on 81 or continues. If it continues infinitely the answer is D

OpenStudy (anonymous):

I think it ends at 81 because it doesnt have "..." i think @wintry

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

I think so too Yes, A

OpenStudy (anonymous):

Do you mind If I discuss a couple more Im not sure about thats it ?

OpenStudy (anonymous):

@wintry

OpenStudy (anonymous):

I don't, post here :-)

OpenStudy (anonymous):

Youre Awesome .. the first is Find the first six terms of the sequence. a1 = -6, an = 4 • an-1 Options are: 0, 4, -24, -20, -16, -12 -24, -96, -384, -1536, -6144, -24,576 -6, -24, -20, -16, -12, -8 -6, -24, -96, -384, -1536, -6144

OpenStudy (anonymous):

the last two are : THANKS :)

OpenStudy (anonymous):

Obviously, it's C or D because of -6 in the first place. a2=-24, a3=-96, so it's D

OpenStudy (anonymous):

thats what I was thinking because i did the -6*4 and what about the next two?

OpenStudy (anonymous):

The second is B. Should I explain why? If you can use a calculator, you can just plug in numbers in formulae and find the right one this way

OpenStudy (anonymous):

the Determine whether the sequence converges or diverges. If it converges, give the limit. one is B? and my calculator broke this week :(

OpenStudy (anonymous):

No, arithmetic sequence, -579+42(n-1)

OpenStudy (anonymous):

what about the one under it ?

OpenStudy (anonymous):

The last one seems to be B

OpenStudy (anonymous):

wait, I need to think a little bit about this one

OpenStudy (anonymous):

okay no problem

OpenStudy (anonymous):

We can find the sequence itself

OpenStudy (anonymous):

It will be \[\frac{ 288 }{ 6^{n}}\]

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

No, why? It's zero, of course

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{ 288 }{ 6^{n} }=0\]

OpenStudy (anonymous):

oh sorry about that that means the answer is B right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are welcome